if a person invests 1000 at an annual rate of 5%, how much interest will the person have after one year
Answer by jim_thompson5910(35256)

Assuming simple interest

I = P*r*t
I = 1000*0.05*1
I = 50

They've earned $50 in interest alone.

They will have 1000+50 = $1,050 in the account after 1 year.